# sixth generation AE autoencoder

### v1

@@
Benedikt Geiger, , Fan Liu, , Shihang Lu, , Andrej Rode, , Daniel Gil Gaviria, , Charlotte Muth, , Laurent Schmalen This work has received funding from the German Federal Ministry of Education and Research (BMBF) within the projects Open6GHub (grant agreement 16KISK010) and KOMSENS-6G (grant agreement 16KISK123). Mr. Geiger acknowledges the Networking Grant from the Karlsruhe House of Young Scientists (KHYS). Parts of this paper have been presented at the IEEE International Symposium on Joint Communications and Sensing (JC&S), Oulu, Finland, January 2025, in [1].Benedikt Geiger, Andrej Rode, Daniel Gil Gaviria, Charlotte Muth, and Laurent Schmalen are with the Communications Engineering Lab (CEL), Karlsruhe Institute of Technology (KIT), Hertzstr. 16, 76187 Karlsruhe, Germany (e-mail: first.lastname@kit.edu). Fan Liu is with the National Mobile Communications Research Laboratory, Southeast University, Nanjing 210096, China (e-mail: fan.liu@seu.edu.cn). Shihang Lu is with the School of Automation and Intelligent Manufacturing, Southern University of Science and Technology, Shenzhen 518055, China (e-mail: lush2021@mail.sustech.edu.cn).
@@ Abstract
Index Terms:
Constellation shaping, Integrated sensing and communication (ISAC), 6G, OFDM, End-to-end learning
@@ I Introduction
Current 5G systems use OFDM due to its numerous and significant advantages, e.g., high spectral efficiency, low-complexity equalization, easy rate adaptation, and easy multi-user downlink handling [4]. Beyond its role in communications, OFDM offers also key advantages for radar systems. In particular, it has the lowest average ranging sidelobe for quadrature amplitude modulation (QAM) constellations [5] among all orthogonal signaling bases, and enables low-complexity range and velocity estimation of targets through fast Fourier transform (FFT)-based processing [6]. This makes OFDM also a compelling choice for future 6G ISAC systems [2].
@@ II System model
In this work, we consider a monostatic OFDM-ISAC system as illustrated in Fig. 1. To reduce computational complexity and simplify the analysis, all sub-carriers use the same constellation, and the sensing target delays are modeled as integer multiples of the sampling period and are shorter than the duration of the cyclic prefix (CP). Furthermore, we assume that the targets are static. We focus on the transmission of single OFDM symbols, and neglect Doppler frequency, i.e., velocity, estimation.
@@ II-B Communications Channel and Receiver
For communications, the influence of a potential multi-path channel is assumed to be mitigated through equalization. The communication receiver removes the CP from the received communication signal $\mathsf{r}_{\text{c}}(t)$ and converts it into the frequency domain using an orthonormal IFFT. This simplifies the OFDM system into a set of $N$ parallel AWGN channels, one for each sub-carrier. Therefore, the received communications symbol of the $n$th sub-carrier after equalization is given by
$\vskip-1.42271pt\mathsf{y}_{\text{c},n}=\mathsf{x}_{n}+\mathsf{w}_{\text{c},n},$ (1)
where $\mathsf{w}_{\text{c},n}\sim\mathcal{CN}(0,\sigma_{\text{c,}n}^{2})$ is AWGN with variance $\sigma_{\text{c,}n}^{2}$. For simplicity, we assume that all sub-carriers observe the same noise variance $\sigma_{\text{c}}^{2}$, i.e., $\sigma_{\text{c,}n}^{2}=\sigma_{\text{c}}^{2},\,\forall n$.
@@
$\vskip-1.42271ptl_{n,m}(y_{\text{c},n})\!=\!\log\frac{\sum_{x_{\!n}\in\mathcal{X}^{(0)}_{m}}{\!f_{\mathsf{y}_{\text{c}}|\mathsf{x}}(y_{\text{c},n}|x_{n})}P_{\mathsf{x}}(x_{n})}{\sum_{x_{n}\in\mathcal{X}^{(1)}_{m}}{\!f_{\mathsf{y}_{\text{c}}|\mathsf{x}}(y_{\text{c},n}|x_{n})}P_{\mathsf{x}}(x_{n})},$ (2)
where $\mathcal{X}^{(b)}_{m}$ represents the set of all constellation symbols labeled with bit $b\in\{0,1\}$ at bit position $m$, and $f_{\mathsf{y}_{\text{c}}|\mathsf{x}}$ denotes the communications channel transition probability density function (PDF).
II-C Achievable Information Rates (AIR)
@@
$\mathbb{I}_{\text{MI}}(\mathsf{x}_{n};\mathsf{y}_{\text{c},n})=\mathbb{E}_{\mathsf{x}_{n},\mathsf{y}_{\text{c},n}}\!\left\{\log_{2}\left(\frac{f_{\mathsf{y}_{\text{c}}|\mathsf{x}}(\mathsf{y}_{\text{c},n}|\mathsf{x}_{n})}{f_{\mathsf{y}_{\text{c}}}(\mathsf{y}_{\text{c},n})}\right)\right\},$ (3)
which defines the highest achievable rate with an arbitrarily complex receiver.
@@
$\mathbb{I}_{\text{GMI}}(\bm{\mathsf{b}}_{n};\mathsf{y}_{\text{c},n})=\left[\mathbb{H}(\bm{\mathsf{b}}_{n})-\sum\limits_{m=1}^{M}\mathbb{H}(\mathsf{b}_{n,m}|\mathsf{y}_{\text{c},n})\right]^{+},$ (4)
@@ II-D Sensing Channel and Receiver
The arbitrary, but fixed sensing channel $h(t)$ is modeled by $J$ static point targets, each with a delay $\tau_{j}$ and a complex amplitude $a_{j}$. The received sensing signal is
$\mathsf{r}_{\text{s}}(t)=h(t)*\mathsf{s}(t)+\mathsf{w}_{\text{s}}(t)=\sum_{j=1}^{J}a_{j}\mathsf{s}(t-\tau_{j})+\mathsf{w}_{\text{s}}(t),\vskip-5.69046pt$ (5)
where $\mathsf{w}_{\text{s}}(t)\!\sim\!\mathcal{CN}\left(0,\sigma_{\text{s}}^{2}\right)$ denotes AWGN with variance $\sigma_{\text{s}}^{2}$. The sensing receiver samples the baseband signal $\mathsf{r}_{\text{s}}(t)$, removes the CP, and transforms it into the frequency domain using the orthonormal FFT. The frequency domain received symbols are
$\mathsf{y}_{\text{s},n}=\mathsf{x}_{n}h_{n}+\mathsf{w}_{\text{s},n}=\mathsf{x}_{n}\frac{1}{\sqrt{N}}\sum_{j=1}^{J}a_{j}\mathrm{e}^{-\mathrm{j}2\pi\frac{n}{N}\tau_{j}}+\mathsf{w}_{\text{s},n},\vskip-5.69046pt$ (6)
@@
$\hat{\mathsf{h}}[k]=\frac{1}{\sqrt{N}}\sum_{n=0}^{N-1}\hat{\mathsf{h}}_{n}\mathrm{e}^{\mathrm{j}2\pi\frac{n}{N}k},\qquad k=\{0,\ldots N-1\}.$ (9)
@@ II-E The Constellation-dependent Detection Probability
We derive the constellation-dependent detection probability by analyzing each signal processing block of the sensing receiver. For the CA-CFAR algorithm with a window length $N_{\text{win}}\rightarrow\infty$, the detection probability converges to
$P_{\text{D}}=P_{\text{FA}}^{\frac{1}{1+\gamma}},\vskip-2.84544pt$ (10)
assuming Gaussian distributed noise and interference [26]. It depends only on the false alarm rate $P_{\text{FA}}$ and the average signal-to-interference-and-noise ratio (SINR) $\gamma$ of the signal fed to the detection algorithm. Since the false alarm rate $P_{\text{FA}}$ is usually fixed during system design, increasing the detection probability $P_{\text{D}}$ requires increasing the average SINR $\gamma$ at the input of the CA-CFAR, i.e., at the output of the IFFT.
@@
The variance
$\sigma^{2}_{\hat{\mathsf{h}}[k]}=\left(\frac{1}{\sqrt{N}}\right)^{\!2}\sum_{n=0}^{N-1}\sigma^{2}_{\hat{\mathsf{h}}_{n}}=\frac{1}{N}N\sigma^{2}_{\hat{\mathsf{h}}_{n}}=\sigma^{2}_{\hat{\mathsf{h}}_{n}}\vskip-5.69046pt$ (11)
of the delay domain channel estimate equals the variance of the frequency domain channel estimate
@@
where $\kappa_{n}$ is the kurtosis of the constellation, which is equivalent to the 4th-order moment for unit power zero mean constellations. Assuming that the targets are independent with uniformly distributed phases, we obtain $|h_{n}|^{2}=\sum_{j=1}^{J}|a_{j}|^{2}$. Moreover, since all subcarriers employ the same constellation, it follows that $\kappa_{n}=\kappa$.
@@
$\frac{1}{\sqrt{N}}\sum_{n=0}^{N-1}\sum_{j=1}^{J}a_{j}\mathrm{e}^{-\mathrm{j}2\pi\frac{n\tau_{j}}{N}}\mathrm{e}^{\mathrm{j}2\pi\frac{nk}{N}}=\begin{cases}\sqrt{N}a_{j},&k\!=\!\tau_{j},\\ 0,&k\!\neq\!\tau_{j}.\end{cases}$ (13)
The fraction of the power in the deterministic (13) and random part (12) yields the average SINR
$\gamma_{\mathrm{TOI}}=\frac{N\cdot|a_{\mathrm{TOI}}|^{2}}{\sum_{j=1}^{J}|a_{j}|^{2}(\kappa-1)+\sigma_{\text{s}}^{2}}$ (14)
of a target of interest (TOI) with complex amplitude $a_{\mathrm{TOI}}$ at the input of the CFAR. From inserting (14) into (10) follows the constellation-dependent detection probability. We observe that the detection probability $P_{\text{D}}$ depends only on the kurtosis $\kappa$ of the constellation for a given scenario.
Therefore, reducing the kurtosis $\kappa$ of the constellation decreases the variance of the channel estimate in the delay and frequency domain, which in turn increases the average SINR and consequently the detection probability $P_{\mathrm{D}}$. Note that unit modulus constellations like phase shift keying (PSK) have the lowest kurtosis $\kappa=$1$\$, leading to the highest detection probability. A circular complex Gaussian PDF, which maximizes the MI for an AWGN channel und an average transmit power constraint, has a kurtosis of $\kappa=$2$\$ and results in a lower detection probability [5].

### v2

@@
Benedikt Geiger, , Fan Liu, , Shihang Lu, , Andrej Rode, , Daniel Gil Gaviria, , Charlotte Muth, , Laurent Schmalen This work has received funding from the German Federal Ministry of Research, Technology and Space (BMFTR) within the projects Open6GHub, Open6GHub+, KOMSENS-6G (grant agreements 16KISK010, 16KIS2405, and 16KISK123). Mr. Geiger acknowledges the Networking Grant from the Karlsruhe House of Young Scientists (KHYS). Parts of this paper have been presented at the IEEE International Symposium on Joint Communications and Sensing (JC&S), Oulu, Finland, January 2025, in [1].Benedikt Geiger, Andrej Rode, Daniel Gil Gaviria, Charlotte Muth, and Laurent Schmalen are with the Communications Engineering Lab (CEL), Karlsruhe Institute of Technology (KIT), Hertzstr. 16, 76187 Karlsruhe, Germany (e-mail: first.lastname@kit.edu). Fan Liu is with the National Mobile Communications Research Laboratory, Southeast University, Nanjing 210096, China (e-mail: fan.liu@seu.edu.cn). Shihang Lu is with the School of Automation and Intelligent Manufacturing, Southern University of Science and Technology, Shenzhen 518055, China (e-mail: lush2021@mail.sustech.edu.cn).
@@ I Introduction
Current 5G systems use OFDM due to its numerous and significant advantages, e.g., high spectral efficiency, low-complexity equalization, easy rate adaptation, and easy multi-user downlink handling [4]. Beyond its role in communications, OFDM offers also key advantages for radar systems. In particular, it has the lowest average ranging sidelobe for quadrature amplitude modulation (QAM) constellations [5] among all orthogonal signaling bases, and enables low-complexity range and velocity estimation of targets through fast Fourier transform (FFT)-based processing [6]. This makes OFDM also a compelling choice for future 6G ISAC systems [2]. Fig. 1 shows a sketch of the ISAC scenario considered in this paper, which follows a standard setup from the literature [2, 7, 8]. The key idea is to use the same transmit signal of, e.g., the base station, for both communications to the user equipment (UE) and sensing. In particular, the ISAC transmit signal is used to convey information to a communication user, and to detect a potential target in a radar-like manner.
Figure 1: Considered ISAC scenario. A base station employs a unified transmit signal with an optimized constellation to communicate with an UE and to sense the environment in a radar-like manner. The objective of the ISAC base station is to transmit data to the UE and to detect a TOI, such as a drone, in the presence of an interfering object like a building. Constellation shaping enables a flexible trade-off between communication throughput and detection probability of potential targets, and it improves both simultaneously compared with legacy modulation formats.
@@ II System model
In this work, we consider a monostatic OFDM-ISAC system as illustrated in Fig. 2. To reduce computational complexity and simplify the analysis, all $N$ sub-carriers use the same constellation, and the sensing target delays are modeled as integer multiples of the sampling period and are shorter than the duration of the cyclic prefix (CP). Furthermore, we assume that the targets are static. We focus on the transmission of single OFDM symbols, and neglect Doppler frequency, i.e., velocity, estimation.
@@ II-B Communications Channel and Receiver
For communications, the influence of a potential multi-path channel is assumed to be mitigated through equalization. The communication receiver removes the CP from the received communication signal $\mathsf{r}_{\text{c}}(t)$ and converts it into the frequency domain using an orthonormal FFT. This simplifies the OFDM system into a set of $N$ parallel AWGN channels, one for each sub-carrier. Therefore, the received communications symbol of the $n$th sub-carrier is given by
$\vskip-1.42271pt\mathsf{y}_{\text{c},n}=\mathsf{x}_{n}+\mathsf{w}_{\text{c},n},\vskip-7.11317pt$ (1)
where $\mathsf{w}_{\text{c},n}\sim\mathcal{CN}(0,\sigma_{\text{c,}n}^{2})$ is AWGN with variance $\sigma_{\text{c,}n}^{2}$.
@@
$\vskip-1.42271ptl_{n,m}(y_{\text{c},n})\!=\!\log\frac{\sum_{x_{\!n}\in\mathcal{X}^{(0)}_{m}}{\!f_{\mathsf{y}_{\text{c,}n}|\mathsf{x}}(y_{\text{c},n}|x_{n})}P_{\mathsf{x}}(x_{n})}{\sum_{x_{n}\in\mathcal{X}^{(1)}_{m}}{\!f_{\mathsf{y}_{\text{c,}n}|\mathsf{x}}(y_{\text{c},n}|x_{n})}P_{\mathsf{x}}(x_{n})},\vskip-7.11317pt$ (2)
where $\mathcal{X}^{(b)}_{m}$ represents the set of all constellation symbols labeled with bit $b\in\{0,1\}$ at bit position $m$, and $f_{\mathsf{y}_{\text{c,}n}|\mathsf{x}}$ denotes the communications channel transition probability density function (PDF).
II-C Achievable Information Rates
@@
$\mathbb{I}_{\text{MI}}(\mathsf{x}_{n};\mathsf{y}_{\text{c},n})=\mathbb{E}_{\mathsf{x}_{n},\mathsf{y}_{\text{c},n}}\!\left\{\log_{2}\left(\frac{f_{\mathsf{y}_{\text{c,}n}|\mathsf{x}}(\mathsf{y}_{\text{c},n}|\mathsf{x}_{n})}{f_{\mathsf{y}_{\text{c,}n}}(\mathsf{y}_{\text{c},n})}\right)\right\},\vskip-7.11317pt$ (3)
which defines the highest achievable rate with an arbitrarily complex receiver, where $f_{\mathsf{y}_{\text{c,}n}}(\mathsf{y}_{\text{c},n})$ denotes the PDF of the received communication symbols $\mathsf{y}_{\text{c,}n}$ on sub-carrier $n$.
@@
$\mathbb{I}_{\text{GMI}}(\bm{\mathsf{b}}_{n};\mathsf{y}_{\text{c},n})=\left[\mathbb{H}(\bm{\mathsf{b}}_{n})-\sum\limits_{m=1}^{M}\mathbb{H}(\mathsf{b}_{n,m}|\mathsf{y}_{\text{c},n})\right]^{+},\vskip-7.11317pt$ (4)
@@ II-D Sensing Channel and Receiver
We consider a sensing scenario with $J$ targets, where one target is designated as the TOI, e.g., a drone, and the remaining J1J-1 targets act as strong interferers, e.g., static buildings or residual self-interference from the finite isolation between the transmitter and receiver paths, which can be modeled as a static target at zero delay with an artificial radar cross section (RCS). The RCS of the TOI follows a Swerling-1 model with PDF
$f_{\mathsf{\sigma}_{\text{RCS,TOI}}}(\sigma_{\text{RCS,TOI}})\!=\!\begin{cases}\!\frac{1}{\bar{\sigma}_{\text{RCS,TOI}}}\exp\!\left(\!\!-\frac{\sigma_{\text{RCS,TOI}}}{\bar{\sigma}_{\text{RCS,TOI}}}\!\right)\!\!,\!\!\!\!&\sigma_{\text{RCS,TOI}}\!\geqslant\!0,\\ \!0,&\sigma_{\text{RCS,TOI}}\!<\!0,\end{cases}\vskip-5.12128pt$ (5)
i.e., the RCS fluctuates independently from OFDM symbol to OFDM symbol according to an exponential distribution with mean $\bar{\sigma}_{\text{RCS,TOI}}$. The interfering targets follow a non-fluctuating Swerling-0 model and therefore have constant RCS values $\sigma_{\mathrm{RCS},j}$, which are larger than the mean RCS of the TOI, i.e., $\sigma_{\mathrm{RCS},j}\gg\bar{\sigma}_{\text{RCS}}$ for all $j\neq\mathrm{TOI}$. This detection setting is widely regarded as particularly challenging, since the sidelobes of the strong interfering targets can obscure the weak TOI. For each OFDM symbol, we draw one realization of the RCS of the TOI and treat it as fixed. Consequently, within a single symbol, the sensing channel can be modeled as the sum of $J$ static point targets with arbitrary, but fixed amplitudes
$\mathsf{r}_{\text{s}}(t)=h(t)*\mathsf{s}(t)+\mathsf{w}_{\text{s}}(t)=\sum_{j=1}^{J}a_{j}\mathsf{s}(t-\tau_{j})+\mathsf{w}_{\text{s}}(t),\vskip-5.69046pt\vskip-2.84544pt$ (6)
where $\mathsf{w}_{\mathrm{s}}(t)\sim\mathcal{CN}(0,\sigma_{\mathrm{s}}^{2})$ denotes AWGN with variance $\sigma_{\text{s}}^{2}$. The amplitudes $a_{j}$ and delays $\tau_{j}$ follow from the radar equation
$a_{j}=\sqrt{\frac{\sigma_{\mathrm{RCS},j}c_{0}^{2}P_{\mathrm{Tx}}G_{\mathrm{Tx}}G_{\mathrm{Rx}}}{(4\pi)^{3}R_{j}^{4}f_{\mathrm{c}}^{2}}}\mathrm{e}^{\mathrm{j}\varphi_{j}},\qquad\tau_{j}=\frac{2R_{j}}{c_{0}},\vskip-5.69046pt$ (7)
where $P_{\mathrm{Tx}}$, $G_{\mathrm{Tx}}$, $G_{\mathrm{Rx}}$, $f_{\mathrm{c}}$, and $c_{0}$ denote transmit power, antenna gains, carrier frequency, and speed of light. The phases $\varphi_{j}$ are independent across targets and, for each target, drawn i.i.d. from a uniform distribution over $0,2\pi)$. The sensing receiver samples the baseband signal $\mathsf{r}_{\text{s}}(t)$, removes the CP, and transforms it into the frequency domain using the orthonormal FFT. The frequency domain received symbols are
$\mathsf{y}_{\text{s},n}=\mathsf{x}_{n}h_{n}+\mathsf{w}_{\text{s},n}=\mathsf{x}_{n}\sum_{j=1}^{J}a_{j}\mathrm{e}^{-\mathrm{j}2\pi\frac{n}{N}\tau_{j}}+\mathsf{w}_{\text{s},n},\vskip-5.12128pt$ (8)
@@
$\hat{\mathsf{h}}[k]=\frac{1}{\sqrt{N}}\sum_{n=0}^{N-1}\hat{\mathsf{h}}_{n}\mathrm{e}^{\mathrm{j}2\pi\frac{n}{N}k},\qquad k=\{0,\ldots N-1\}.\vskip-4.26773pt$ (11)
@@ II-E The Constellation-dependent Detection Probability
We derive the constellation-dependent detection probability of the TOI by analyzing each signal processing block of the sensing receiver. For the CA-CFAR algorithm with a window length $N_{\text{win}}\rightarrow\infty$, the detection probability is
$P_{\text{D}}=P_{\text{FA}}^{\frac{1}{1+\bar{\gamma}}_{\text{TOI}}},\vskip-2.84544pt$ (12)
assuming Gaussian distributed noise and interference [26]. The detection probability $P_{\text{D}}$ depends only on the false alarm rate $P_{\text{FA}}$ and the average signal-to-interference-and-noise ratio (SINR) $\bar{\gamma}_{\text{TOI}}$ at the input of the detector. To prevent clutter-induced bursts of false alarms and to enable robust and predictable detection performance, the false alarm rate $P_{\text{FA}}$ is usually fixed during system design [26]. Consequently, increasing the detection probability $P_{\text{D}}$ requires increasing the average SINR $\bar{\gamma}_{\text{TOI}}$ at the input of the CA-CFAR, i.e., at the output of the IFFT. Note that the average is taken with respect to the RCS fluctuations of the TOI, as modeled by the Swerling-1 model. Following standard practice [26], we first derive the SINR for a fixed sensing-channel realization (6) and subsequently average over the sensing-channel fluctuations.
@@
Since the frequency-domain channel estimation errors are zero-mean and uncorrelated on different subcarriers, and using the linearity of the FFT, the variance
$\sigma^{2}_{\hat{\mathsf{h}}[k]}=\left(\frac{1}{\sqrt{N}}\right)^{\!2}\sum_{n=0}^{N-1}\sigma^{2}_{\hat{\mathsf{h}}_{n}}$ (13)
of the delay domain channel estimate equals the mean variance of the frequency domain channel estimates
@@
where $\kappa_{n}$ is the kurtosis of the constellation
$\kappa_{n}=\frac{\mathbb{E}_{\mathsf{x}}\!\left\{\left|\mathsf{x}_{n}-\mathbb{E}_{\mathsf{x}}\!\left\{\mathsf{x}_{n}\right\}\right|^{4}\right\}}{\left(\mathbb{E}_{\mathsf{x}}\!\left\{\left|\mathsf{x}_{n}-\mathbb{E}_{\mathsf{x}}\!\left\{\mathsf{x}_{n}\right\}\right|^{2}\right\}\right)^{2}}\underset{\mathbb{E}_{\mathsf{x}}\!\left\{|\mathsf{x}_{n}|^{2}\right\}=1}{\overset{\mathbb{E}_{\mathsf{x}}\!\left\{\mathsf{x}_{n}\right\}=0}{=}}\mathbb{E}_{\mathsf{x}}\!\left\{|\mathsf{x}_{n}|^{4}\right\},\vskip-1.42271pt$ (15)
which is equivalent to the 4th-order moment for unit power zero mean constellations. Moreover, since all subcarriers employ the same constellation, it follows that $\kappa_{n}=\kappa$. Inserting (14) into (13) yields the noise-and-interference power
$\sigma^{2}_{\hat{\mathsf{h}}[k]}=\sum_{j=1}^{J}|a_{j}|^{2}(\kappa-1)+\sigma_{\text{s}}^{2},\vskip-5.69046pt$ (16)
where we exploit that all targets are independent with uniformly distributed phases
$\frac{1}{N}\sum_{n=1}^{N-1}|h_{n}|^{2}=\sum_{j=1}^{J}|a_{j}|^{2}.\vskip-8.5359pt$ (17)
@@
$\frac{1}{\sqrt{N}}\sum_{n=0}^{N-1}\sum_{j=1}^{J}a_{j}\mathrm{e}^{-\mathrm{j}2\pi\frac{n\tau_{j}}{N}}\mathrm{e}^{\mathrm{j}2\pi\frac{nk}{N}}=\begin{cases}\sqrt{N}a_{j},&k\!=\!\tau_{j},\\ 0,&k\!\neq\!\tau_{j}.\end{cases}\vskip-1.42271pt$ (18)
The fraction of the power in the deterministic (18) and random part (14) yields the SINR
$\gamma_{\mathrm{TOI}}=\frac{N\cdot|a_{\mathrm{TOI}}|^{2}}{\sum_{j=1}^{J}|a_{j}|^{2}(\kappa-1)+\sigma_{\text{s}}^{2}}\vskip-2.84544pt$ (19)
for a fixed sensing channel realization at the input of the CFAR algorithm.
TABLE I: Kurtosis values of modulation formats
Constellation $\widetilde{M}-$PSK $16$-QAM $64$-QAM $256$-QAM $\mathcal{CN}(0,1)$
Kurtosis $\kappa$ 11 $1.320$ $1.381$ $1.395$ 22
We observe that the SINR depends only on the kurtosis $\kappa$ of the constellation for a given sensing channel realization. Using the identity $\mathrm{Var}\{|\mathsf{x}|^{2}\}=\mathbb{E}\{|\mathsf{x}|^{4}\}-|\mathbb{E}\{|\mathsf{x}|^{2}\}|^{2}$ and the unit-power constraint $\mathbb{E}\{|\mathsf{x}|^{2}\}=1$, the kurtosis can be written as
$\kappa=\mathrm{Var}\{|\mathsf{x}|^{2}\}+1,\vskip-2.84544pt$ (20)
showing that the kurtosis directly measures the variance of the amplitudes of the constellation around the unit circle. Table I lists the kurtosis of common modulation formats. Unit-modulus constellations such as phase shift keying (PSK) achieve the minimum kurtosis $\kappa=1$, preventing interference since the interference term in (16) vanishes and only the AWGN noise power $\sigma_{\mathrm{s}}^{2}$ remains. In contrast, conventional modulation formats such as $\widetilde{M}-$QAM and a circular complex Gaussian PDF, which maximizes the MI for an AWGN channel under an average transmit power constraint, introduce interference. For conventional $\widetilde{M}-$QAM, the kurtosis approaches $\kappa\approx 1.4$ as $M\to\infty$ (see Appendix A), while a circular complex Gaussian PDF has $\kappa=$2$\$.
We observe that the interference introduced by non-constant-modulus constellations, i.e., $\kappa>1$, scales with the sum of the powers of all targets. Since the TOI is assumed to be much weaker than the interfering targets, the interference power is dominated by the non-fluctuating interferers, such that the resulting noise-and-interference term is effectively independent of the RCS fluctuations of the TOI. In contrast, the received target power itself fluctuates according to the Swerling-1 model. Inserting (7) into (19) and averaging with respect to these fluctuations yields the average SINR of the TOI
$\bar{\gamma}_{\text{TOI}}=\frac{N\cdot|\bar{a}_{\mathrm{TOI}}|^{2}}{\sum_{j=1}^{J}|a_{j}|^{2}(\kappa-1)+\sigma_{\text{s}}^{2}},$ (21)
where $\bar{a}_{\mathrm{TOI}}$ denotes the average amplitude corresponding to the mean RCS $\bar{\sigma}_{\text{RCS,TOI}}$. Substituting (21) into (12) gives the constellation-dependent detection probability. Consequently, the RCS fluctuations of the TOI average out, and the detection probability depends only on the mean amplitude of the TOI, the amplitude of the interfering targets, and the kurtosis $\kappa$ of the constellation.